# An object-centric core metamodel for IoT-enhanced event logs

### v1

@@ 3.4.3 CAIRO
The CAIRO metamodel, shown in Figure 5, has been proposed in (Franceschetti et al. 2023c), motivated by the need for an IoT-driven process monitoring and conformance checking across a wide range of processes and domains. The CAIRO metamodel aims at being agnostic with respect to process and domain characteristics such as level of process structuredness, process modelling paradigm (e.g., imperative or declarative), level of automation of the process activities, and process-awareness, i.e., the support from a Process-aware information system for orchestrating process executions. The foundational assumptions are the observability of process events in terms of IoT events, the knowledge of the monitored process (either formally or informally specified), and a defined mapping of low-level IoT events into high-level process events. At the core of the metamodel is the notion of CAIRO event, i.e., an event that is contextualized, Atomic, Instantaneous, Relevant, and Observable. The contextualized property is represented in the metamodel by the association between the CAIRO event class and the Event attribute class, which posits that event attributes associate contextual information to an event. The property of being atomic (i.e., events have all-or-nothing semantics) is enforced in the metamodel by cardinality constraints associating each instance of an Event occurrence of a CAIRO event with exactly one instance of a Process event from a known process specification. The property of being instantaneous binds a CAIRO event to a specific point in time, and is realized by associating a CAIRO event with exactly one timestamp attribute. The property of being relevant requires an event to refer to a relevant state change of a process, and is enforced in the metamodel by associating the CAIRO event class with the Monitoring point class, which models applications detecting activity occurrences based on IoT events. The property of being observable is realized in the CAIRO metamodel with the association between class Monitoring point and class Expression, representing the aggregation of the readings from instances of the associated class Sensor.
Figure 5: CAIRO metamodel (Franceschetti et al. 2023c)
@@ 4 Methodology
The proposed CORE metamodel is the artifact resulting from the application of the design science research methodology described in (Hevner et al. 2004). According to this methodology, the development of the artifact represented as the CORE metamodel follows a sequence of activities. These activities are based on the Design Science Research Methodology (DSRM) process model by (Peffers et al. 2008), which consists of six activities: 1) Identify Problem and Motivate, 2) Define Objectives of a Solution, 3) Design and Development of an Artifact, 4) Demonstration, 5) Evaluation, and 6) Communication. This manuscript serves as the medium of Communication of the developed artifact. The remaining activities are described in the following sections, along with insights on how they have been conducted in the context of the development of the CORE metamodel .
@@ 4.1 Identify Problem and Motivate
The development of a common metamodel is motivated by the need to integrate IoT data with process event data in IoT-enhanced business processes–a task complicated by the heterogeneity, noise, and low granularity of IoT sensor data. As shown in Sect. 3, current models differ significantly in their assumptions, application domains, and design objectives, resulting in fragmentation that impedes interoperability and data reuse. A common metamodel aims to bridge these gaps by capturing core concepts shared across existing models, facilitating translation and standardization of logs across diverse use cases. This promotes broader adoption by fostering the exchange of IoT-enhanced process data.
@@ 4.2 Define Objectives of a Solution
To develop a suitable research artifact, concrete objectives for the solution have to be defined. Here, the objective is to provide a common metamodel for the representation of IoT sensor data and business process data in a unified way. For this purpose, requirements have been derived from multiple research groups that work in the field at the intersection of IoT and BPM. These research groups have already developed their own models to represent IoT-enriched event logs based on specific requirements. An initial set of nine requirements has been collected from the already available models by analyzing their respective requirements, identifying commonalities, and discussing priorities for a common metamodel. Afterward, this set of requirements has been refined through several discussion rounds, finally resulting in seven requirements that are the basis for the development of the CORE metamodel .
@@
    – Different Granularity Levels: Different granularity levels must be representable, including a) IoT events (e.g., sensor data) and b) process events (e.g., status changes in workflows).
@@
    – Traceability: Traceability across different granularity levels is essential. Therefore, it should be possible a) to establish relationships between low-level and high-level events, such as linking IoT events to process events. Furthermore, since event abstraction can be part of functions or algorithms, b) it should be possible to annotate an event with the abstraction algorithm used, indicating that it was generated through this process (de Murillas et al. 2019).
@@
    – Semantic Annotations: Prior knowledge should be usable if leveraged. Therefore, the log should be able to represent a) procedural knowledge, such as information about the process model (formal or informal, e.g., in natural language), b) the relationship between sensors and the process within the infrastructure, and c) prior knowledge from other data sources, such as databases or taxonomies. This semantic knowledge provides sensor relationships (e.g., two sensors measuring different dimensions of the same feature, two sensors being dependent on each other, …), or about collected data (e.g., how should it behave generally, how should it be visualized, …).
@@
    – Flexible Case Notion: For some use-cases, a single case notion is not sufficient. Therefore, a more flexible case notion is desired, which allows to not only link events to one specific case but to (potentially multiple) entities being involved in its execution. This enables analysis from different points of view (i.e., based on different corresponding entities such as machines, locations, …) (de Murillas et al. 2019).
@@
    – Different Data Sources: Data in an IoT environment can originate from different sources, and it should be possible to integrate and represent various data sources, including a) information systems (IS), such as ERP systems or databases containing process metadata, b) sensors, which provide real-time data like temperature, humidity, or motion, and c) actuators, which record actions such as opening a valve, starting a motor, or adjusting a device’s settings (de Murillas et al. 2019).
@@
    – Represent Metadata:
    The log should also include metadata that describe the broader context or configuration in which events are recorded. Unlike semantic annotations, metadata applies to the log or logging setup as a whole rather than individual events. Typical examples include: a) Information about the data acquisition environment (e.g., machine configurations, sensor firmware versions, log generation settings); b) Timestamps or durations relating to log generation periods (e.g., start/end of logging, time zone offsets); and c) Global properties relevant to interpretation (e.g., process version, organizational unit, or logging strategy). This metadata is essential for ensuring correct interpretation, comparison, and reuse of event logs across systems and settings.
@@
    To highlight the need for a common metamodel, Table 2, provides an overview of the fulfillment of the above requirements by the models presented in Section 3.4. It is clear from the table that each existing model has varying degrees of fulfillment of the requirements, which align with the specific goals driving their development. For instance, NICE is designed to be very flexible, which translates into fulfilling R2 and R5, while CAIRO focuses on knowledge-intensive processes, as evidenced by its fulfillment of R4 and R7.
@@ 4.3 Design and Development of an Artifact
The development of the CORE metamodel required a structured and scientifically grounded alignment process among international experts from six specialized working groups. To achieve this, a participatory and iterative development approach has been adopted, drawing on the principles of Participatory Design (Bødker et al. 2022) and iterative prototyping (Camburn et al. 2017). This methodology ensured the active participation of all stakeholders while allowing continuous validation and refinement of the model.
@@ 5.1 Main concepts derivation from existing models
To create the common CORE metamodel , we started by analyzing existing models to 1) identify similarities and 2) distinguish features that addressed the common requirements (necessary features for the common model; in some cases decide which solution to keep) from features that are specific to the models’ use cases (e.g. ambiguity in CAIRO) or features that violate the requirements (e.g. single case notion in DS)
@@
Then, in order to 1) reuse tooling and 2) ensure visibility, we integrated these concepts within the OCEL framework. The main reason for linking our model with OCEL is its growing popularity in the process mining community, among both researchers and practitioners, which we think will be crucial to enhance the visibility of our model. Moreover, this allows us to reuse the broad software tooling readily available to manipulate and analyse object-centric event data. Concretely, this means that we translated the common core concepts in either events and objects, the two main concepts in the OCEL metamodel.
Concretely, we refined and specialized the concepts of event and object with subtypes to precisely define the events and objects used in IoT-enhanced event logging.
Finally, we derived relationships between the different subtypes of objects and events that we defined. To do so, we started from the OCEL metamodel, which defines the two following relationships:
@@
Note that there is no direct event-event relationship; events can only be related indirectly, via a common object. To obtain the relationships in the CORE metamodel , we adapted the relationships in existing models to respect theses constraints from the OCEL metamodel.
In the remainder of this section, we first present a high-level view of our metamodel, before diving deeper into how we modeled the concepts retained from the original IoT-enhanced log models. Then, we give a concrete example of use of the CORE metamodel from the running use case.
@@ 5.2 Metamodel presentation
Note that, in addition to objects and events, we also include Time, to remain consistent with OCEL 2.0.
@@
  • Business objects: object of interest for the process, which are impacted by events. They can correspond to case notions, or they can define the scope of a context parameter;
  • General objects: other types of objects, which do not fall into one of the two previously introduced classes, are explicitly mentioned for emphasis. Examples include subprocesses or activity objects (Benzin et al. 2024).
Each event needs exactly one data source and one or more business objects, and can be linked to objects of other classes (but that is not mandatory).
@@ 5.2.1 Event type definitions
In process mining, an event is commonly defined as the execution of (part of) an activity Günther and Verbeek (2014). However, an event has a very different meaning in the IoT literature, where it is often seen as an observation made by a sensor Janowicz et al. (2019); Bertrand et al. (2024). To resolve this conflict, our model introduces two types of events, IoT events and Process events, and gives them explicit definitions.
An IoT event represents a change in a physical phenomenon, which is typically derived from one observation of one IoT device such as a sensor, e.g., a measurement made by a temperature sensor. To account for complex physical phenomena, which cannot directly be tracked by an IoT device, but require to process the observations made by one IoT device or to aggregate data collected by multiple IoT devices, we also make it possible to derive one higher-level IoT event from one or more lower-level IoT events Seiger et al. (2023), which can be characterized by a label indicating the meaning of an event. An example of such an IoT event is the detection of a peak in time series sensor data, which requires processing consecutive observations of one sensor using an algorithm.
@@ 5.2.2 Objects taxonomy
In our metamodel, we distinguish two potentially overlapping types of business objects: Case objects and Context objects.
First, the Case objects represent traditional objects as defined by the OCED (Fahland et al. 2024) and OCEL standards (Berti et al. 2024). Such objects are entities that are involved in the execution of activities. As such, they represent cases in subprocesses that provide different views on long end-to-end processes. Examples include orders, product bottles, batches, customers, etc.
@@ 6 Proof-Of-Concept Implementation
This section outlines the Python implementation of CORE metamodel , outlining the architectural decisions and software components developed. It serves as a proof-of-concept [^3] https://github.com/chimenkamp/core-model-for-event-logs and a practical contribution, enabling researchers and practitioners to apply CORE metamodel to their specific use cases. First, we provide an overview of the architecture and motivate our design decisions (Section 6.1). Furthermore, we describe the input layer, where the interfaces and parsers are defined (Section 6.2). Next, we report on the main logic, contained in the data transformation layer (Section 6.3). Finally, we illustrate how the data can be represented using the OCEL standard, ensuring that our approach remains independent of any particular OCEL implementation (Section 6.4).
@@ 6.1 Architectural Overview
The decision of an architecture for CORE metamodel adheres to specific requirements (e.g., model conformity, rapid prototyping). Model conformity ensures correctness and consistency. Rapid prototyping allows quick iterations due to frequent metamodel changes. Given the fast and dynamic CORE metamodel development, we prioritize simplicity, reliability, and testability as key requirements for the architecture decision. Based on these requirements and following established best practices (Richards and Ford 2020), we decided to use a Layered Software Architecture as an architectural style for our implementation. Following this, we define three layers: the Input Layer is necessary to parse the different existing logs (see Section 3.4). The Data Transformation Layer links concepts between the input model and CORE metamodel (see Section 5.1 and Figure 2). Finally, the OCEL Representation Layer transforms and represents our model in OCEL (see Figure 6). The architecture is visualized in Figure 10, consisting of the three layers, Input Layer, Data Transformation Layer, OCEL Representation Layer. To closely reflect our implementation, we explicitly include Python-specific type annotations in the UML representations.
@@ 6.2 Input Layer
The Input Layer addresses the need for compatibility with the diverse set of existing models that CORE metamodel must support (see Section 3.4). This compatibility ensures that the transformation process begins with a correct and semantically aligned interpretation of heterogeneous event formats. It consists of two main components (Figure 11). First, the abstract parser, handles the translation of existing IoT-enhanced event log formats (e.g., NICE, CAIRO, DataStream) with pre-defined classes. In addition, it can be utilized to write custom parsers. This component defines a common interface for parsing different log formats into the internal data structures. To illustrate the concept transfer, we consider the DataStream parser as an example in the following. The DataStream format utilizes two different types of events: process events that represent high-level activities (e.g., “Start Production”, “Quality Check”), and stream points that capture IoT sensor observations. For each process event, the parser creates corresponding objects (e.g., machines, resources) and establishes relationships between them. For instance, when a machine operator starts a production process, this creates a process event with an associated resource object (the operator). Simultaneously, stream points are translated into observation objects and linked to the corresponding process events. Any values not directly required by CORE metamodel (e.g., ‘system type’, ‘interaction type’) are preserved as attributes.

### v2

@@ 3.4.3 CAIRO
The CAIRO metamodel, shown in Figure 5, has been proposed in (franceschetti2023event), motivated by the need for an IoT-driven process monitoring and conformance checking across a wide range of processes and domains. The CAIRO metamodel aims at being agnostic with respect to process and domain characteristics such as level of process structuredness, process modelling paradigm (e.g., imperative or declarative), level of automation of the process activities, and process-awareness, i.e., the support from a Process-aware information system for orchestrating process executions. The foundational assumptions are the observability of process events in terms of IoT events, the knowledge of the monitored process (either formally or informally specified), and a defined mapping of low-level IoT events into high-level process events. At the core of the metamodel is the notion of CAIRO event, i.e., an event that is contextualized, Atomic, Instantaneous, Relevant, and Observable. The contextualized property is represented in the metamodel by the association between the CAIRO event class and the Event attribute class, which posits that event attributes associate contextual information to an event. The property of being atomic (i.e., events have all-or-nothing semantics) is enforced in the metamodel by cardinality constraints associating each instance of an Event occurrence of a CAIRO event with exactly one instance of a Process event from a known process specification. The property of being instantaneous binds a CAIRO event to a specific point in time, and is realized by associating a CAIRO event with exactly one timestamp attribute. The property of being relevant requires an event to refer to a relevant state change of a process, and is enforced in the metamodel by associating the CAIRO event class with the Monitoring point class, which models applications detecting activity occurrences based on IoT events. The property of being observable is realized in the CAIRO metamodel with the association between class Monitoring point and class Expression, representing the aggregation of the readings from instances of the associated class Sensor.
Figure 5: CAIRO metamodel (franceschetti2023event)
@@ 4 Methodology
The proposed CORE metamodel is the artifact resulting from the application of the design science research methodology described in (Hevner.2004_DSRinIS). According to this methodology, the development of the artifact represented as the CORE metamodel follows a sequence of activities. These activities are based on the Design Science Research Methodology (DSRM) process model by (Peffers.2008_DSRM_ProcessModel), which consists of six activities: 1) Identify Problem and Motivate, 2) Define Objectives of a Solution, 3) Design and Development of an Artifact, 4) Demonstration, 5) Evaluation, and 6) Communication. This manuscript serves as the medium of Communication of the developed artifact. The remaining activities are described in the following sections, along with insights on how they have been conducted in the context of the development of the CORE metamodel .
@@ 4.1 Identify Problem and Motivate
The development of a common metamodel is motivated by the need to integrate IoT data with process event data in IoT-enhanced business processes–a task complicated by the heterogeneity, noise, and low granularity of IoT sensor data (bertrand2022assessing). As shown in Sect. 3, current models differ significantly in their assumptions, application domains, and design objectives, resulting in fragmentation that impedes interoperability and data reuse. A common metamodel aims to bridge these gaps by capturing core concepts shared across existing models, facilitating translation and standardization of logs across diverse use cases. This promotes broader adoption by fostering the exchange of IoT-enhanced process data.
@@ 4.2 Define Objectives of a Solution
To develop a suitable research artifact, concrete objectives for the solution have to be defined. Here, the objective is to provide a common metamodel for the representation of IoT sensor data and business process data in a unified way. For this purpose, requirements have been derived from multiple research groups that work in the field at the intersection of IoT and BPM. These research groups have already developed their own models to represent IoT-enhanced event logs based on specific requirements. An initial set of requirements has been collected from the already available models by analyzing their respective requirements, identifying commonalities, and discussing priorities for a common metamodel. Afterward, this set of requirements has been refined through several discussion rounds, finally resulting in seven requirements that are the basis for the development of the CORE metamodel .
@@
    – Different Granularity Levels: Different granularity levels must be representable, including a) IoT events (e.g., low-level sensor data, higher level IoT events (seiger2023data)) and b) process events (e.g., state changes in workflow executions).
@@
    – Traceability: Traceability across different granularity levels is essential. Therefore, it should be possible a) to establish relationships between low-level and high-level events, such as linking IoT events to process events. Furthermore, since event abstraction can be part of functions or algorithms, b) it should be possible to annotate an event with the abstraction algorithm or expression used, indicating that it was generated through this process (GonzalezLopezdeMurillas2019).
@@
    – Semantic Annotations: In semantics-aware process analytics (rebmann2022enabling), prior knowledge can be used, if available. Therefore, the metamodel should be able to represent a) procedural knowledge, such as information about the process model (formal or informal, e.g., in natural language), b) the relationship between sensors and the process within the infrastructure, and c) prior knowledge from other data sources, such as databases or taxonomies.
@@
    – Flexible Case Notion: For some use-cases, a single case notion is not sufficient. Therefore, a more flexible case notion is desired, which allows to not only link events to one specific case but to (potentially multiple) entities being involved in its execution. This enables analysis from different points of view (i.e., based on different corresponding entities such as machines, locations, etc. (GonzalezLopezdeMurillas2019).
@@
    – Different Data Sources: Data in an IoT environment can originate from different sources, which are very heterogeneous. It should be possible to integrate and represent various data sources, including a) information systems (IS), such as ERP systems or databases containing process metadata, b) sensors, which provide real-time data like temperature, humidity, or motion, and c) actuators, which record actions such as opening a valve, starting a motor, or adjusting a device’s settings (GonzalezLopezdeMurillas2019). Moreover, it should be possible to represent these sources in a unified manner, despite the high level of heterogeneity.
@@
    – Represent Metadata: The log should also include metadata that describe the broader context or configuration in which events are recorded. Unlike semantic annotations, metadata applies to the log or logging setup as a whole rather than individual events. Typical examples include: a) Information about the data acquisition environment (e.g., machine configurations, sensor firmware versions, log generation settings); b) Timestamps or durations relating to log generation periods (e.g., start/end of logging, time zone offsets); and c) Global properties relevant to interpretation (e.g., logging strategy). This metadata is essential for ensuring correct interpretation, comparison, and reuse of event logs across systems and settings.
@@
    To highlight the need for a common metamodel, Table 2, provides an overview of the fulfillment of the above requirements by the models presented in Section 3.4. It is clear from the table that each existing model has varying degrees of fulfillment of the requirements, which align with the specific goals driving their development. For instance, NICE is designed to be very flexible, which translates into fulfilling R2 and R5, while CAIRO focuses on knowledge-intensive processes, as evidenced by its fulfillment of R4 and R7 via event attribute, monitoring point, and process specification.
@@ 4.3 Design and Development of an Artifact
The development of the CORE metamodel required a structured and scientifically grounded alignment process among international experts from six specialized working groups. To achieve this, a participatory and iterative development approach has been adopted, drawing on the principles of Participatory Design (bodker2022participatory) and iterative prototyping (camburn2017design). This methodology ensured the active participation of all stakeholders while allowing continuous validation and refinement of the model.
@@ 5.1 Main Concepts Derivation from Existing Models
To create the common CORE metamodel , we started by analyzing existing models to 1) identify similarities and 2) distinguish features that addressed the common requirements (necessary features for the common model; in some cases decide which solution to keep) from features that are specific to the models’ use cases or features that violate the requirements (e.g., single case notion in DataStream).
@@
Then, in order to 1) reuse tooling and 2) ensure visibility, we integrated these concepts within the OCEL 2.0 metamodel. The main reason for linking our model with OCEL 2.0 is its growing popularity in the process mining community, among both researchers and practitioners, which we think will be crucial to enhance the visibility and uptake of our model. Moreover, this allows us to reuse the broad software tooling readily available to manipulate and analyze object-centric event data. Concretely, this means that we translated the common core concepts in either events and objects, the two main concepts in the OCEL 2.0 metamodel.
We therefore refined event and object into explicit subtypes to precisely define the events and objects used in IoT-enhanced event logging.
Finally, we derived relationships between the different subtypes of objects and events that we defined. To do so, we started from the OCEL 2.0 metamodel, which defines the two following relationships:
@@
Note that there is no direct event-event relationship; events can only be related indirectly, via a common object. To obtain the relationships in the CORE metamodel , we adapted the relationships in existing models to respect theses constraints from the OCEL 2.0 metamodel.
In the remainder of this section, we first present a high-level view of our metamodel, before diving deeper into how we modeled the concepts retained from the original IoT-enhanced log models. Then, we give a concrete example of use of the CORE metamodel from the running example.
@@ 5.2 Metamodel Presentation
Note that, in addition to objects and events, we also include Time for consistency with OCEL 2.0.
@@
  • Business objects: object of interest for the process, which are impacted by events. They can correspond to process models, cases, or they can define the scope of a context parameter;
  • General objects: other types of objects, which do not fall into one of the two previously introduced classes, are explicitly mentioned for emphasis. Examples include subprocesses or activity objects (benzin2024inexa).
Each event needs exactly one data source and one or more business objects, and can (but does not have to) be linked to objects of other classes.
@@ 5.2.1 Event Type Definitions
In process mining, an event is commonly defined as the execution of (part of) an activity gunther2014xes. However, an event has a very different meaning in the IoT literature, where it is often seen as an observation made by a sensor janowicz2019sosa; Bertrand2024Feb. To resolve this discrepancy, our model introduces two types of events, IoT events and process events, and gives them explicit definitions.
An IoT event represents a change in a physical phenomenon, which is typically derived from one observation of one IoT device such as a sensor, e.g., a measurement made by a temperature sensor. To account for complex physical phenomena, which cannot directly be tracked by an IoT device, but require to process the observations made by one IoT device or to aggregate data collected by multiple IoT devices, we also make it possible to derive a higher-level IoT event from one or more lower-level IoT events seiger2023data, which can be characterized by a label indicating the semantics of the event. An example of such an IoT event is the detection of a peak in time series sensor data, which requires processing consecutive observations of one sensor using an algorithm.
@@ 5.2.2 Objects Taxonomy
In our metamodel, we distinguish between two potentially overlapping types of business objects: Case objects and Context objects.
First, the Case objects represent traditional objects as defined by the OCED (fahland2024towards) and OCEL standards (berti2024ocel). Such objects are entities that are involved in the execution of activities. As such, they represent cases in subprocesses that provide different views on long end-to-end processes. Examples include orders, product bottles, batches, etc.
@@ 6 Proof-Of-Concept Implementation
This section outlines the Python implementation of CORE metamodel , outlining the architectural decisions and software components developed. It serves as a proof-of-concept [^4] https://github.com/chimenkamp/core-model-for-event-logs and a practical contribution, enabling researchers and practitioners to apply CORE metamodel to their specific use cases. First, we provide an overview of the architecture and motivate our design decisions (Section˜6.1). Furthermore, we describe the input layer, where the interfaces and parsers are defined (Section˜6.2). Next, we report on the main logic, contained in the data transformation layer (Section˜6.3). Finally, we illustrate how the data can be represented using the OCEL standard, ensuring that our approach remains independent of any particular OCEL implementation (Section˜6.4).
@@ 6.1 Architectural Overview
The decision of an architecture for CORE metamodel adheres to specific requirements (e.g., model conformity, rapid prototyping). Model conformity ensures correctness and consistency. Model-driven engineering (MDE) allows quick iterations due to frequent metamodel changes. Given the dynamic CORE metamodel development, we prioritize simplicity, reliability, and testability as key requirements for the architecture decision. Based on these requirements and following established best practices (ford2020fundamentals), we decided to use a Layered Software Architecture as an architectural style for our implementation. Following this, we define three layers: the Input Layer is necessary to parse the different existing logs (see Section˜3.4). The Data Transformation Layer links concepts between the input model and CORE metamodel (see Section˜5.1 and Figure˜2). Finally, the OCEL Representation Layer transforms and represents our model in OCEL (see Figure˜6). The architecture is visualized in Figure˜10, consisting of the three layers, Input Layer, Data Transformation Layer, OCEL Representation Layer. To closely reflect our implementation, we explicitly include Python-specific type annotations in the UML representations.
@@ 6.2 Input Layer
The Input Layer addresses the need for compatibility with the diverse set of existing models that CORE metamodel must support (see Section˜3.4). This compatibility ensures that the transformation process begins with a correct and semantically aligned interpretation of heterogeneous event formats. It consists of two main components (Figure˜11). First, the abstract parser, handles the translation of existing IoT-enhanced event log formats (e.g., NICE, CAIRO, DataStream) with pre-defined classes. For that, the user will need to map the the components (i.e., from the custom format) to the CORE metamodel components. In addition, it can be utilized to write custom parsers. This component defines a common interface for parsing different log formats into the internal data structures. To illustrate the concept transfer, we consider the DataStream parser as an example in the following. The DataStream format utilizes two different types of events: process events that represent high-level activities (e.g., “Start Production”, “Quality Check”), and stream points that capture IoT sensor observations. For each process event, the parser creates corresponding objects (e.g., machines, resources) and establishes relationships between them. For instance, when a machine operator starts a production process, this creates a process event with an associated resource object (the operator). Simultaneously, stream points are translated into observation objects and linked to the corresponding process events. Any values not directly required by CORE metamodel (e.g., ‘system type’, ‘interaction type’) are preserved as attributes.